# StatLLaMA: A multi-stage training framework for building a domain-optimized statistical language model

### v1

@@ 2.6 Reinforcement learning from human feedback
While SFT and instruction tuning enable LLMs to perform tasks and follow natural language instructions, they are insufficient for ensuring alignment with complex human values, ethical standards, and preferences. To address this gap, reinforcement learning from human feedback (RLHF) has become a widely adopted and highly successful alignment technique. Instead of manually designing reward functions, RLHF derives reward signals indirectly from human preference comparisons over model-generated outputs.
@@
    Reward model training: Human annotators rank or select preferred responses among multiple model outputs $\{y_{1},y_{2},\cdots,y_{n}\}$ for a given prompt $x$. These rankings are converted into pairwise comparisons $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$. A reward model $\text{RM}_{\phi}$ is trained to assign higher scores to preferred responses using a Bradley–Terry–based objective [Taorietal2023alpacaa]:
@@
Proximal policy optimization (PPO) [Schulmanetal2017proximalpolicy, Zhengetal2023secretsof] is the most widely used reinforcement learning algorithm in RLHF, valued for its stability and efficiency. It employs importance sampling to reuse data generated by an old policy and introduces a clipping mechanism that constrains policy updates, preventing large deviations and stabilizing training. This balance of sample efficiency and stability makes PPO particularly well-suited for large language model alignment. Despite its success in aligning LLMs [Devlinetal2019bertpre-training, Radfordetal2018improvinglanguage, Touvronetal2023llamaopen], PPO remains complex, resource-intensive, and sensitive to hyperparameters. Challenges include designing robust reward models, preventing the pattern collapse (instability in policy learning) [Zhengetal2023secretsof], and mitigating the alignment tax (reduced performance on traditional NLP tasks) [Linetal2024mitigatingthe].
To address these issues, alternative preference optimization methods have been proposed. Direct preference optimization (DPO) [Rafailovetal2023directpreference] simplifies training by bypassing the reward model and learning directly from preference pairs with a classification-style objective. It derives its loss function from the KL-regularized RLHF objective and uses a closed-form reparameterization based on the Bradley-Terry model. Advantages of DPO include: no need for a reward model or value network, greater training stability, and lower computational cost. However, DPO may be sensitive to the quality of the SFT model and can exhibit asymmetric learning dynamics when preference margins are small [Pan2025etalwhatmatters].
Group relative policy optimization (GRPO) [Shaoetal2024deepseekmathpushing], a lightweight PPO variant, eliminates the value network and estimates the advantage function using group-average reward baselines. It retains PPO’s clipping mechanism and KL regularization, providing stable policy updates with reduced memory consumption. GRPO is particularly suited for applications with limited compute resources, high-throughput preference training, and step-by-step supervision needs.
@@ 2.7 Knowledge distillation
Knowledge distillation (KD) is a model compression technique designed to transfer knowledge from a large, high-performing teacher model (T) to a smaller, more efficient student model (S) [Hintonetal2015distillingthe]. The objective is to reduce inference costs without significantly sacrificing performance, making models more suitable for deployment on resource-limited or latency-sensitive systems.
In natural language generation, KD can be applied at different levels. For word-level KD (logits KD), the student is trained to match the teacher’s probability distribution over the next token at each time step by minimizing the KL divergence between them [Calderonetal2023asystematic]. This allows the student to capture fine-grained linguistic and semantic knowledge but requires direct access to the teacher’s logits—often infeasible when the teacher is a closed-source API. For sequence-level KD, instead of imitating token-level distributions, the student learns from full output sequences $\hat{y}_{T}$ generated by the teacher for inputs $x$. These sequences act as pseudo-targets (PTs) [Calderonetal2023asystematic, Shleiferetal2020pre-trainedsummarization]. This approach is more practical in black-box settings since it requires only the teacher’s final outputs, not intermediate logits.
In this study, we adopt sequence-level KD. Using the teacher model’s API, we generated a large set of pseudo-labeled pairs $(x,\hat{y}_{T})$, which serve as training data for SFT on the student model. The objective minimizes the NLL Loss (Eq. 1) of the student when predicting PTs. The main advantage of this method is its simplicity and practicality: it requires only teacher outputs, not internal states or logits. Furthermore, pseudo-labeled data from KD has proven valuable in reinforcement learning alignment pipelines, providing supervision signals for reward model training or policy optimization. This enables the construction of competitive preference models without costly human annotation, making KD a powerful complement to RLHF.
@@ 2.8 Downstream task fine-tuning
Even after RLHF alignment, model preferences often remain biased toward general-purpose language generation. To optimize performance for specific applications, a final stage of downstream task fine-tuning (DTFT) is applied. This process uses a relatively small amount of task-specific labeled data to make fine adjustments to the aligned model, ensuring that outputs better satisfy domain requirements and evaluation criteria.
Recent studies [Hsiehetal2023distillingstep-by-step] highlight that DTFT not only improves task adherence but can also transfer more complex capabilities, thereby enhancing the adaptability of smaller models in specialized, task-oriented applications. In practice, sequence-level knowledge distillation often serves as an effective initialization or data augmentation strategy to quickly build a student model with basic language understanding and generation capabilities, but its full potential is realized only when combined with DTFT.
@@ 3.2.2 Instruction tuning
3.2.3 Supervised fine-tuning (SFT)
@@
3.2.4 Reinforcement learning from human feedback (RLHF)
@@
3.2.5 Downstream task fine-tuning (DTFT)
@@ 3.3.1 Grade School Math 8K (GSM8K)
GSM8K, released by OpenAI, is a widely used benchmark for assessing multi-step mathematical reasoning. It contains approximately 8,500 natural language math word problems, each paired with a complete step-by-step solution. This study adopted the 8-shot CoT evaluation paradigm (GSM8K (Test)), in which the model receives eight annotated examples of “question–reasoning–answer” before being asked a new question. This setup encourages the model to emulate step-by-step logical reasoning during inference. The primary evaluation metric is answer accuracy, reflecting the model’s capability in logical reasoning and numerical computation.
@@ 4.1.2 Processing of PDF data
PDF data constituted a large portion of our corpora, including Statistics Open-Source Books, Statistics Course Books, and arXiv (stat.*). Due to their unstructured nature, we designed a specialized workflow. Text and block coordinates were extracted using PyMuPDF (fitz) [McKie2024pymupdfa], forming the basis for semantic reconstruction. We also experimented with combining title keywords and block positions to identify chapter boundaries, particularly in arXiv papers. However, due to significant variability in document layouts, this method could not consistently and accurately segment chapters and was therefore excluded from the main pipeline.
For multimodal processing, we explored using the BLIP model [Lietal2022blipbootstrapping] to generate textual descriptions of statistical charts to supplement image content. Yet, because of the model’s limited accuracy in interpreting professional charts and the difficulty of reliably matching figures with annotations, these descriptions were not incorporated into the core workflow. For mathematical formulas, we evaluated Pix2Text [Breezedeus2024pix2textan] for converting image-based equations into LaTeX. While reasonably effective for simple formulas, it struggled with complex layouts and introduced speed and compatibility issues. We therefore retained formulas in their original text style, allowing the model to infer their meaning from surrounding context.
Finally, to mitigate duplication caused by page breaks or formatting artifacts, we implemented a similarity detection mechanism based on Levenshtein distance [Levenshtein1965binarycodes], which identified and removed repeated content in adjacent paragraphs, thereby improving corpus purity and training efficiency.
@@ 4.1.3 Processing of structured sources
Beyond plain text, we incorporated structured knowledge sources, such as the STATO [STATO2025] and OBCS [Zhengetal2016theontology] ontologies. These provide explicit semantics and relationships in a knowledge graph (KG) format. Using pandas [reback2020pandas], we cleaned the raw data and built a mapping dictionary linking ontology entity URIs to their preferred labels. We then iterated through triplets (subject, predicate, object) to generate standardized text statements. To improve relational coverage, inverse triplets were automatically added following predefined rules. Duplicates were removed via set operations, yielding a clean collection of structured statements suitable for model learning and integration into stages such as SFT.
@@ 4.1.4 Extracting structured knowledge from text
As a supplementary exploration, we attempted to extract structured semantic knowledge directly from unstructured academic text. Using SpaCy’s en_core_web_lg model [Honnibaletal2020spacyindustrial-strength] with dependency parsing, we extracted subject-verb-object (SVO) triplets. In principle, this approach could support automatic construction of a semantic network. However, the complexity of statistical writing—long sentences, subordinate clauses, and technical terminology—led to low accuracy and limited coverage compared to curated ontologies. As such, these results were excluded from the main training corpus and retained only for auxiliary analysis.
@@ 4.1.5 Unified formatting and tokenization
Once cleaning and structural modeling were complete, all text was converted into numerical sequences for training. We adopted a pre-tokenization strategy, ensuring all text was tokenized and ID-mapped prior to training, which reduced overhead and improved efficiency. Tokenization was performed using Hugging Face’s AutoTokenizer [AutoTokenizer2020], configured to match the vocabulary and algorithm of the base model. Texts exceeding the model’s maximum sequence length were processed using a sliding window approach to preserve context, while shorter inputs were padded. Padding tokens were assigned a label value of –100, ensuring they did not affect loss computation.
@@ 4.2.1 Foundation models and parameter-efficient fine-tuning
This study adopts Meta’s LLaMA-3.2-3B as the base foundation model for all fine-tuning and experimental procedures. The model is built on an optimized Transformer decoder-only architecture, featuring several key improvements over the original Transformer design [Ibe2024unlockinglow-resource]. LLaMA-3.2-3B was selected for this study due to its open availability, competitive performance, and excellent compatibility with parameter-efficient fine-tuning (PEFT) methods.
PEFT in this study primarily employed LoRA, which introduces low-rank trainable matrices into targeted model layers, enabling efficient adaptation with minimal parameter updates. LoRA’s rank and scaling factor ($\alpha$) were adjusted dynamically based on task type and training stage. To further reduce memory consumption and computation costs, the base model was loaded using 4-bit quantization via the bitsandbytes library [Dettmersetal2023bitsandbytes]. Training employed the AdamW optimizer, with gradient accumulation used to achieve an effective batch size appropriate for available hardware resources.
@@ 4.2.2 Training and fine-tuning framework
For large-scale training during the CoP stage or for distributed multi-GPU fine-tuning, we integrated the DeepSpeed framework [Rasley2020deepspeedsystem]. Developed by Microsoft, DeepSpeed implements advanced distributed optimization strategies, most notably the ZeRO (zero redundancy optimizer), which shards model parameters, gradients, and optimizer states across multiple GPUs, thereby minimizing memory overhead. This study utilized ZeRO Stage 2 with 8 A100 GPUs provided by Taiwan Computing Cloud, significantly accelerating training throughput. While DeepSpeed offers excellent scalability, integration required careful configuration and compatibility tuning with the Hugging Face Trainer, along with additional engineering considerations such as loss monitoring and runtime stability.
@@ 5.3 Pipeline 3: An efficient path for domain adaptation and optimization
5.3.2 RLHF preference optimization: From GRPO’s limitations to the successful application of DPO
@@
GRPO experiments: As an initial attempt along the RLHF-style preference optimization path, we applied group relative policy optimization (GRPO) to the SFT-v1 and SFT-v2 models. For SFT-v2, we further explored three GRPO configurations by varying: LoRA rank and alpha, and the data mixing ratio between Statistical GRPO and GSM8K (Train) samples. Table 6 summarizes the GRPO settings, and Table 7 reports the benchmark scores.
@@
DPO Experiments: We therefore shifted our focus to direct preference optimization (DPO), which reframes preference learning as a direct classification-style optimization on paired preference data, avoiding explicit reward modeling and complex reinforcement learning updates. Theoretically, this offers a more stable and straightforward training process.
@@ 5.3.3 Downstream task fine-tuning: Effects of different fine-tuning intensities
Having obtained SFT-v3.4 + DPO as a strong and well-balanced statistical LLM, Pipeline 3 proceeds to its final stage: downstream task fine-tuning (DTFT). This stage serves two core purposes:
@@ 6.1 Summary of findings
This study systematically examined multi-stage training strategies for developing a statistical large language model (LLM) capable of integrating domain-specific statistical knowledge with strong general-purpose reasoning abilities. Using the lightweight LLaMA-3.2-3B family as the foundation, our objective was to construct a model deployable in resource-constrained environments while achieving high accuracy on tasks requiring statistical expertise.
By empirically comparing three representative training pipelines, we demonstrate that the choice of starting model critically determines the efficiency and success of domain adaptation. Pipelines that begin from the base model—without pre-existing instruction-following capability—struggle to absorb domain knowledge, even after extensive continual pretraining (CoP) and subsequent instruction tuning. As a result, performance on core statistical reasoning tasks remains limited, and later supervised fine-tuning (SFT) or preference-alignment stages often yield marginal or even negative gains.
@@
In the alignment stage, direct preference optimization (DPO) consistently outperformed group relative policy optimization (GRPO). GRPO demonstrated high sensitivity to hyperparameters and produced unstable capability trade-offs across reasoning metrics. DPO, by contrast, improved output quality reliably, aligned responses more closely with disciplinary norms in statistics, and—importantly—restored general abilities diminished during SFT. The best-performing model from the SFT stage (v3.4) improved further under DPO, producing the most balanced and robust model in the study.
Finally, our analysis of downstream task fine-tuning (DTFT) revealed that when a model has undergone substantial prior optimization, only extremely low-intensity fine-tuning is viable. Conventional or moderate-intensity fine-tuning quickly destabilizes earlier gains and induces substantial performance degradation. A carefully calibrated low-step, low-rank LoRA configuration (DTFT-v2) successfully adapted the model to high-quality domain-specific QA data while preserving core capabilities. Even slight increases in training intensity led to disproportionate declines. Additional experiments incorporating statistical conversation data produced mixed benefits, suggesting the importance of careful alignment between data type and evaluation objectives.
@@
High-quality Instruct initialization $\rightarrow$ Domain-oriented SFT $\rightarrow$ Stable preference alignment (DPO) $\rightarrow$ Minimal, controlled DTFT.

### v2

@@ 2.6 Reinforcement learning from human feedback
While SFT and instruction tuning enable LLMs to perform tasks and follow natural language instructions, they are insufficient for ensuring alignment with complex human values, ethical standards, and preferences. To address this gap, RLHF has become a widely adopted and highly successful alignment technique. Instead of manually designing reward functions, RLHF derives reward signals indirectly from human preference comparisons over model-generated outputs.
@@
    Reward model training: Human annotators rank or select preferred responses among multiple model outputs $\{y_{1},y_{2},\cdots,y_{n}\}$ for a given prompt $x$. These rankings are converted into pairwise comparisons $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$. A reward model $\text{RM}_{\phi}$ is trained to assign higher scores to preferred responses using a Bradley–Terry–based objective [60]:
@@
PPO [54, 70] is the most widely used reinforcement learning algorithm in RLHF, valued for its stability and efficiency. It employs importance sampling to reuse data generated by an old policy and introduces a clipping mechanism that constrains policy updates, preventing large deviations and stabilizing training. This balance of sample efficiency and stability makes PPO particularly well-suited for large language model alignment. Despite its success in aligning LLMs [15, 49, 63], PPO remains complex, resource-intensive, and sensitive to hyperparameters. Challenges include designing robust reward models, preventing the pattern collapse (instability in policy learning) [70], and mitigating the alignment tax (reduced performance on traditional NLP tasks) [40].
To address these issues, alternative preference optimization methods have been proposed. DPO [51] simplifies training by bypassing the reward model and learning directly from preference pairs with a classification-style objective. It derives its loss function from the KL-regularized RLHF objective and uses a closed-form reparameterization based on the Bradley-Terry model. Advantages of DPO include: no need for a reward model or value network, greater training stability, and lower computational cost. However, DPO may be sensitive to the quality of the SFT model and can exhibit asymmetric learning dynamics when preference margins are small [47].
GRPO [55], a lightweight PPO variant, eliminates the value network and estimates the advantage function using group-average reward baselines. It retains PPO’s clipping mechanism and KL regularization, providing stable policy updates with reduced memory consumption. GRPO is particularly suited for applications with limited compute resources, high-throughput preference training, and step-by-step supervision needs.
@@ 2.7 Knowledge distillation
Knowledge distillation (KD) is a model compression technique designed to transfer knowledge from a large, high-performing teacher model (T) to a smaller, more efficient student model (S) [21]. The objective is to reduce inference costs without significantly sacrificing performance, making models more suitable for deployment on resource-limited or latency-sensitive systems.
In natural language generation, KD can be applied at different levels. For word-level KD (logits KD), the student is trained to match the teacher’s probability distribution over the next token at each time step by minimizing the KL divergence between them [8]. This allows the student to capture fine-grained linguistic and semantic knowledge but requires direct access to the teacher’s logits—often infeasible when the teacher is a closed-source API. For sequence-level KD, instead of imitating token-level distributions, the student learns from full output sequences $\hat{y}_{T}$ generated by the teacher for inputs $x$. These sequences act as pseudo-targets [8, 57]. This approach is more practical in black-box settings since it requires only the teacher’s final outputs, not intermediate logits.
In this study, we adopt sequence-level KD. Using the teacher model’s API, we generated a large set of pseudo-labeled pairs $(x,\hat{y}_{T})$, which serve as training data for SFT on the student model. The objective minimizes the NLL Loss (Eq. 1) of the student when predicting pseudo-targets. The main advantage of this method is its simplicity and practicality: it requires only teacher outputs, not internal states or logits. Furthermore, pseudo-labeled data from KD has proven valuable in reinforcement learning alignment pipelines, providing supervision signals for reward model training or policy optimization. This enables the construction of competitive preference models without costly human annotation, making KD a powerful complement to RLHF.
@@ 2.8 Downstream task fine-tuning
Even after RLHF alignment, model preferences often remain biased toward general-purpose language generation. To optimize performance for specific applications, a final stage of DTFT is applied. This process uses a relatively small amount of task-specific labeled data to make fine adjustments to the aligned model, ensuring that outputs better satisfy domain requirements and evaluation criteria.
Recent studies [25] highlight that DTFT not only improves task adherence but can also transfer more complex capabilities, thereby enhancing the adaptability of smaller models in specialized, task-oriented applications. In practice, sequence-level knowledge distillation often serves as an effective initialization or data augmentation strategy to quickly build a student model with basic language understanding and generation capabilities, but its full potential is realized only when combined with DTFT.
@@ 3.2.2 Instruction tuning
3.2.3 Supervised fine-tuning
@@
3.2.4 Reinforcement learning from human feedback
@@
3.2.5 Downstream task fine-tuning
@@ 3.3.1 Grade School Math 8K (GSM8K)
GSM8K, released by OpenAI, is a widely used benchmark for assessing multi-step mathematical reasoning. It contains approximately 8,500 natural language math word problems, each paired with a complete step-by-step solution. This study adopted the 8-shot CoT evaluation paradigm (GSM8K (Test)), in which the model receives eight annotated examples of “question-reasoning-answer” before being asked a new question. This setup encourages the model to emulate step-by-step logical reasoning during inference. The primary evaluation metric is answer accuracy, reflecting the model’s capability in logical reasoning and numerical computation.
@@ 4.1.2 Processing of PDF data
PDF data constituted a large portion of our corpora, including Statistics Open-Source Books, Statistics Course Books, and arXiv (stat.*). Due to their unstructured nature, we designed a specialized workflow. Text and block coordinates were extracted using PyMuPDF (fitz) [45], forming the basis for semantic reconstruction. We also experimented with combining title keywords and block positions to identify chapter boundaries, particularly in arXiv papers. However, due to significant variability in document layouts, this method could not consistently and accurately segment chapters and was therefore excluded from the main pipeline.
For multimodal processing, we explored using the BLIP model [37] to generate textual descriptions of statistical charts to supplement image content. Yet, because of the model’s limited accuracy in interpreting professional charts and the difficulty of reliably matching figures with annotations, these descriptions were not incorporated into the core workflow. For mathematical formulas, we evaluated Pix2Text [6] for converting image-based equations into LaTeX. While reasonably effective for simple formulas, it struggled with complex layouts and introduced speed and compatibility issues. We therefore retained formulas in their original text style, allowing the model to infer their meaning from surrounding context.
Finally, to mitigate duplication caused by page breaks or formatting artifacts, we implemented a similarity detection mechanism based on Levenshtein distance [35], which identified and removed repeated content in adjacent paragraphs, thereby improving corpus purity and training efficiency.
@@ 4.1.3 Processing of structured sources
Beyond plain text, we incorporated structured knowledge sources, such as the STATO [62] and OBCS [69] ontologies. These provide explicit semantics and relationships in a knowledge graph (KG) format. Using pandas [61], we cleaned the raw data and built a mapping dictionary linking ontology entity URIs to their preferred labels. We then iterated through triplets (subject, predicate, object) to generate standardized text statements. To improve relational coverage, inverse triplets were automatically added following predefined rules. Duplicates were removed via set operations, yielding a clean collection of structured statements suitable for model learning and integration into stages such as SFT.
@@ 4.1.4 Extracting structured knowledge from text
As a supplementary exploration, we attempted to extract structured semantic knowledge directly from unstructured academic text. Using SpaCy’s en_core_web_lg model [22] with dependency parsing, we extracted subject-verb-object (SVO) triplets. In principle, this approach could support automatic construction of a semantic network. However, the complexity of statistical writing—long sentences, subordinate clauses, and technical terminology—led to low accuracy and limited coverage compared to curated ontologies. As such, these results were excluded from the main training corpus and retained only for auxiliary analysis.
@@ 4.1.5 Unified formatting and tokenization
Once cleaning and structural modeling were complete, all text was converted into numerical sequences for training. We adopted a pre-tokenization strategy, ensuring all text was tokenized and ID-mapped prior to training, which reduced overhead and improved efficiency. Tokenization was performed using Hugging Face’s AutoTokenizer [28], configured to match the vocabulary and algorithm of the base model. Texts exceeding the model’s maximum sequence length were processed using a sliding window approach to preserve context, while shorter inputs were padded. Padding tokens were assigned a label value of –100, ensuring they did not affect loss computation.
@@ 4.2.1 Foundation models and parameter-efficient fine-tuning
This study adopts Meta’s LLaMA-3.2-3B as the base foundation model for all fine-tuning and experimental procedures. The model is built on an optimized Transformer decoder-only architecture, featuring several key improvements over the original Transformer design [29]. LLaMA-3.2-3B was selected for this study due to its open availability, competitive performance, and excellent compatibility with PEFT methods.
PEFT in this study primarily employed LoRA, which introduces low-rank trainable matrices into targeted model layers, enabling efficient adaptation with minimal parameter updates. LoRA’s rank and scaling factor ($\alpha$) were adjusted dynamically based on task type and training stage. To further reduce memory consumption and computation costs, the base model was loaded using 4-bit quantization via the bitsandbytes library [13]. Training employed the AdamW optimizer, with gradient accumulation used to achieve an effective batch size appropriate for available hardware resources.
@@ 4.2.2 Training and fine-tuning framework
For large-scale training during the CoP stage or for distributed multi-GPU fine-tuning, we integrated the DeepSpeed framework [53]. Developed by Microsoft, DeepSpeed implements advanced distributed optimization strategies, most notably the ZeRO (zero redundancy optimizer), which shards model parameters, gradients, and optimizer states across multiple GPUs, thereby minimizing memory overhead. This study utilized ZeRO Stage 2 with 8 A100 GPUs provided by Taiwan Computing Cloud, significantly accelerating training throughput. While DeepSpeed offers excellent scalability, integration required careful configuration and compatibility tuning with the Hugging Face Trainer, along with additional engineering considerations such as loss monitoring and runtime stability.
@@ 5.3 Pipeline 3: An efficient path for domain adaptation and optimization
5.3.2 RLHF preference optimization: From GRPO limitations to effective DPO alignment
@@
GRPO experiments: As an initial attempt along the RLHF-style preference optimization path, we applied GRPO to the SFT-v1 and SFT-v2 models. For SFT-v2, we further explored three GRPO configurations by varying: LoRA rank and alpha, and the data mixing ratio between Statistical GRPO and GSM8K (Train) samples. Table 6 summarizes the GRPO settings, and Table 7 reports the benchmark scores.
@@
DPO Experiments: We therefore shifted our focus to DPO, which reframes preference learning as a direct classification-style optimization on paired preference data, avoiding explicit reward modeling and complex reinforcement learning updates. Theoretically, this offers a more stable and straightforward training process.
@@ 5.3.3 Downstream task fine-tuning: Effects of different fine-tuning intensities
Having obtained SFT-v3.4 + DPO as a strong and well-balanced statistical LLM, Pipeline 3 proceeds to its final stage: DTFT. This stage serves two core purposes:
@@ 6.1 Summary of findings
This study systematically examined multi-stage training strategies for developing a statistical LLM capable of integrating domain-specific statistical knowledge with strong general-purpose reasoning abilities. Using the lightweight LLaMA-3.2-3B family as the foundation, our objective was to construct a model deplorable in resource-constrained environments while achieving high accuracy on tasks requiring statistical expertise.
By empirically comparing three representative training pipelines, we demonstrate that the choice of starting model critically determines the efficiency and success of domain adaptation. Pipelines that begin from the base model—without pre-existing instruction-following capability—struggle to absorb domain knowledge, even after extensive CoP and subsequent instruction tuning. As a result, performance on core statistical reasoning tasks remains limited, and later SFT or preference-alignment stages often yield marginal or even negative gains.
@@
In the alignment stage, DPO consistently outperformed GRPO. GRPO demonstrated high sensitivity to hyperparameters and produced unstable capability trade-offs across reasoning metrics. DPO, by contrast, improved output quality reliably, aligned responses more closely with disciplinary norms in statistics, and—importantly—restored general abilities diminished during SFT. The best-performing model from the SFT stage (v3.4) improved further under DPO, producing the most balanced and robust model in the study.
Finally, our analysis of DTFT revealed that when a model has undergone substantial prior optimization, only extremely low-intensity fine-tuning is viable. Conventional or moderate-intensity fine-tuning quickly destabilizes earlier gains and induces substantial performance degradation. A carefully calibrated low-step, low-rank LoRA configuration (DTFT-v2) successfully adapted the model to high-quality domain-specific QA data while preserving core capabilities. Even slight increases in training intensity led to disproportionate declines. Additional experiments incorporating statistical conversation data produced mixed benefits, suggesting the importance of careful alignment between data type and evaluation objectives.
@@
High-quality Instruct initialization $\rightarrow$ Domain-oriented SFT $\rightarrow$ Stable DPO preference alignment $\rightarrow$ Minimal, controlled DTFT.